40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 20 21:43:25 2010, yuta, Update, CDS, checking whitening filters for MCs and messed up MC2ULSEN.pngMC2LRSEN.png
    Reply  Thu Oct 21 21:35:16 2010, yuta, Update, CDS, fixed whitening filter switches LLLR.pngContec23BO.png
    Reply  Fri Oct 22 00:59:01 2010, yuta, Update, CDS, new MEDM screens, new filter banks(whitening and dewhitening related) HappySideMEDM.png
       Reply  Fri Oct 22 23:29:48 2010, yuta, Update, CDS, fixed output filter switching 
          Reply  Mon Nov 1 23:31:05 2010, yuta, Update, CDS, checked input hardware filter in single frequency 
Message ID: 3767     Entry time: Fri Oct 22 23:29:48 2010     In reply to: 3758     Reply to this: 3842
Author: yuta 
Type: Update 
Category: CDS 
Subject: fixed output filter switching 

(Joe, Yuta)

Summary:
 This evening, we fixed output filter analog/digital switchings in Simulink logic, but they didn't actually switched filters.
 That was because what we thought BO0 was BO2 and BO2 was BO0.
 We re-wired, re-labeled the cables.
 We checked it by using a LED(see "test configuration" in this wiki page).
 Also, we checked the Simulink connection by probing MAX333A in SOS Dewhitening Board located at 1X4-8-11B, which switches SIDE coils for all optics.

Next work:

 - Currently, FM10 in UL/UR/LR/LL/SDCOIL filters switch analog/digital, but let's make FM9 do the switching.(See schematic in elog #3758)
 - Check all the logic by measuring transfer functions(maybe single frequency measurement is enough. we can use tdsdmd etc to automate).

ELOG V3.1.3-